OpenStudy (anonymous):

Write an expression that represents the area of the triangle with the base 2x to the power of 2 and height 3xy ( for any triangle with base b and height h, A= 1/2 bh. )

mathslover (mathslover):

area of a triangle = 1/2 * base * height \[\huge{2x^2 * 3xy / 2}\] \[\huge{\frac{6x^3y}{2}=\frac{3x^3y}{1}}\]

mathslover (mathslover):

\[\huge{3x^3y}\]

OpenStudy (anonymous):

thank u <333